# FLAMINGO: combining kinetic SZ effect and galaxy-galaxy lensing measurements to gauge the impact of feedback on large-scale structure

### v1

@@ Abstract
pubyear: 2024pagerange: FLAMINGO: combining kinetic SZ effect and galaxy-galaxy lensing measurements to gauge the impact of feedback on large-scale structureB
@@ 1 Introduction
Various possible solutions have been put forward to potentially reconcile the low-redshift LSS observations with the primary CMB and BAO data. These include unidentified or mischaracterised systematic uncertainties in the LSS observations, or possibly even in the primary CMB measurements. On the theoretical side, LSS tests of cosmology often probe deep into the non-linear regime, and it has been proposed that new physics on those scales could explain the tension (e.g., Amon & Efstathiou 2022; Preston et al. 2023), such as new dark sector models (e.g. Rogers et al., 2023; Elbers et al., 2024), or a mis-understanding of galaxy formation effects (e.g. McCarthy et al., 2018). Modelling non-linear scales necessitates cosmological simulations (or models calibrated on such simulations) to predict the clustering of matter on small scales and at late times. Additionally, as matter collapses to form self-gravitating halos, densities increase to the point where radiative cooling of the gas becomes efficient, leading to further collapse and galaxy formation. This process is accompanied by various energetic feedback mechanisms related to star formation and the accretion of matter onto supermassive black holes.
@@ 2.1 FLAMINGO simulations
The FLAMINGO suite consists of 16 hydrodynamical simulations presented in Schaye et al. (2023), two decaying dark matter variants in Elbers et al. (2024), plus two new hydrodynamical simulations introduced here (see below), and 12 gravity-only simulations. The suite has variations in resolution, box size, subgrid modelling, and cosmology. We use the intermediate resolution ($m_{\mathrm{gas}}=1.09\times 10^{9}$ M) in box sizes of (1 Gpc)3. These simulations use $2\times 1800^{3}$ gas and dark matter particles and $1000^{3}$ neutrino particles, and most adopt cosmological parameters corresponding to the maximum likelihood DES Y3 ‘3×2pt + All Ext.’ $\Lambda$CDM cosmology (Abbott et al., 2022), which we refer to as ‘D3A’. These values assume a spatially flat universe and are based on the combination of constraints from DES Y3 ‘$3\times 2$-point’ correlation functions: cosmic shear, galaxy clustering, and galaxy-galaxy lensing, with constraints from external data from BAO, redshift-space distortions, SN Type Ia, and Planck observations of the CMB (including CMB lensing), Big Bang nucleosynthesis, and local measurements of the Hubble constant (see Abbott et al. 2022 for details). We also consider two alternative cosmologies: a run with the Planck 2018 maximum likelihood cosmology (‘Planck’; Planck Collaboration et al. 2020) and the ‘lensing cosmology’ from Amon et al. (2023) (‘LS8’). The latter model has a lower amplitude of the power spectrum, $S_{8}=0.766$, compared with 0.815 and 0.833 for D3A and Planck, respectively.
The simulations were run with the cosmological smoothed particle hydrodynamics and gravity code SWIFT (Schaller et al., 2024) using the SPHENIX SPH scheme (Borrow et al., 2022). The initial conditions are obtained from a modified version of monofonIC (Hahn et al., 2021; Elbers et al., 2022), and neutrinos are implemented with the $\delta f$ method (Elbers et al., 2021). The modelling of subgrid physics (star formation, stellar evolution, radiative cooling, and sources of feedback) is described in Schaye et al. (2023) and references therein.
@@ 2.1.1 Lightcones: HEALPix maps and galaxy/halo catalogs
The HEALPix map resolution is set to $N_{\text{side}}=16384$, which gives a maximum pixel radius of 13.46 arcseconds and $12*16384^{2}$ pixels in each full sky map. We note that the number of pixels exceeds the size of a signed 32-bit integer ($2^{31}$), which would prevent us from smoothing the kSZ effect maps with the ACT beam (necessary for a like-with-like comparison) using the Healpy smoothing function (sphtfunc.smoothing), as the function can currently only handle a maximum $N_{\text{side}}$ of 8192. We therefore downsample the kSZ effect maps to this resolution. We describe the production of galaxy-galaxy lensing and kSZ effect maps in Sections 3.2 and 3.3, respectively.
Computational limitations prevent us from running a structure finder on-the-fly during the simulation. Instead, to produce galaxy/halo catalogs, structure finding [^1] We have checked that our results and conclusions are not strongly dependent on the choice of substructure finder, by comparing our fiducial results with those derived from employing an independent halo catalog derived using the VELOCIraptor package (Elahi et al., 2019). is done in post-processing on the snapshot particle data using a modified version of the HBT+ algorithm (Han et al., 2012, 2018, Forouhar Moreno et al. in prep). Note that at $z\leq 3$ snapshots are also written out with a frequency of $\Delta z=0.05$, which was chosen to minimise any issues arising from the evolution of galaxy/halo properties between the snapshots and the HEALPix maps. We read in the subhalo catalog corresponding to a given snapshot and then read in a spherical shell from the black hole (BH) particle lightcone which spans the redshift range that is half way to the previous snapshot to halfway to the next snapshot. Every time a BH particle in the lightcone shell appears as the most bound BH in a subhalo in the snapshot (identified by matching their unique particle IDs), we place a copy of the subhalo at the BH particle’s position [^2] Note that subhaloes can potentially contain multiple black holes. To address this issue, we identify which of the BHs in a subhalo exist at both the next and previous snapshots and we pick the most bound BH particle from that subset to use as a tracer. In cases where the most bound BH survives we then use the same BH as before. If the most bound BH is swallowed before the next snapshot, we pick the next most bound BH in the same subhalo (which is not about to get swallowed) and use that as the tracer. in the lightcone. We repeat this process for every snapshot to make the full lightcone halo catalog.
@@ 3.2 Galaxy-galaxy lensing profiles
As noted previously, we downsample our full resolution $N_{\text{side}}=16384$ kSZ effect maps to $N_{\text{side}}=8192$ so that we can then smooth the maps with a 1.3 arcmin Gaussian beam for comparison with the ACT 150 GHz measurements.
@@ 4.1 Dependence on feedback and cosmology
Figure 2: Comparison of the 150 GHz Planck+ACT kSZ effect temperature profiles of BOSS LOWZ and CMASS galaxies from Schaan et al. (2021) (data points with $1\sigma$ error bars) with predictions from the fiducial FLAMINGO run (L1_m9) and the fgas feedback variants. The bottom x-axis shows the angular scale (in arcmins) while the top x-axis shows the comoving projected radius at $z=0.31$ and $z=0.54$ for the LOWZ and CMASS panels, respectively. The vertical dashed lines indicate the location of $R_{\textrm{500c}}$ derived from the mean halo mass of each redshift bin (see Section 4.3). The solid curves represent the predictions for the best-fitting minimum stellar mass (fitted to the lensing, see Fig. 1). The legend provides the number of standard deviations that the model deviates from the observational measurements with the error bars corresponding to the propagated uncertainties on the stellar mass cut given the uncertainties on the galaxy-galaxy lensing measurements. Note that the number of standard deviations is computed using the full covariance matrices of the data, which is important since the outermost bins are highly correlated. The error bars on the data points reflect only the diagonal elements of the covariance matrices. The bottom sub-panels in each of the main panels show the ratio of the fgas feedback variants and the observational data with respect to the fiducial model. For reference, the dotted curves in the right panel show the predicted $\Delta T_{\rm kSZ}$ profiles for the fiducial FLAMINGO run with minimum stellar mass $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]$ values ranging between 10.9 and 11.6 in steps of 0.1. The fiducial calibrated FLAMINGO run is statistically ruled out by the kSZ effect measurements at about the $7\sigma$ level, for both the LOWZ and CMASS samples which are independent. Only the strongest feedback model in this comparison (fgas-$8\sigma$) is formally consistent with the measurements, with a $\approx 1\sigma$ deviation from the measurements.
The solid curves show the best-fitting profiles for each of the FLAMINGO runs, which fall nearly on top of each other. The legend provides the best-fitting ($\log_{10}$) minimum stellar mass (in solar masses) of the simulated galaxy selection for each of the runs. For example, in the LOWZ-L1 bin, the FLAMINGO fiducial feedback run with a minimum stellar mass of $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.24\pm 0.03$ provides the best fit to the DES Y3 + KiDS 1000 $\Delta\Sigma$ measurements. The error bars reflect the $2\sigma$ uncertainties (95% confidence) on the best-fitting minimum mass given the uncertainties on the lensing measurements. The best-fitting minimum stellar mass and its uncertainties are estimated by first calculating the $\Delta\Sigma$ profiles at different values of the minimum stellar mass and then interpolating to obtain the best fit result. The bottom right panel of Fig. 1 shows the process, where the dotted curves correspond to the predicted $\Delta\Sigma$ profiles for the fiducial FLAMINGO run with $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]$ values ranging between 10.9 and 11.6 in steps of 0.1. We then interpolate the $\Delta\Sigma$ values to a much finer grid of $\log_{10}{M_{\text{star}}}$ in bins of projected radius, $R$. The best-fitting minimum stellar mass is determined by a $\chi^{2}$ minimisation assuming that the radial bins are uncorrelated.
@@
Comparing the fits across the two LOWZ bins (top two panels), for the majority of the cases the preferred minimum stellar masses are consistent within a few sigma for a given run. For example, the fiducial FLAMINGO model prefers a minimum stellar mass of $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.24\pm 0.03$ for the LOWZ-L1 bin and $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.29\pm 0.02$ for the LOWZ-L2 bin. As the two LOWZ bins are consistent with a single LOWZ selection for a given model, we jointly fit the LOWZ L1 and L2 bins to determine the minimum stellar mass cut for the kSZ effect predictions. The same is true for the higher redshift CMASS bins (C1 and C2, with $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.14\pm 0.05$ and $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.23\pm 0.04$, respectively), although we note that for the CMASS-C1 bin the shape of the best-fitting simulated profiles do not match the measurements perfectly. In particular, we note that the 1-halo regime dominates the fit (in terms of signal to noise), and the simulations provide a good match to the data there, but they predict a signal that is too large compared to the measurements at larger radii, in the 2-halo regime. The CMASS-C1 bin was also identified as an outlier in Amon et al. (2023), who found that even their flexible HOD framework (with 5 free parameters) was unable to reproduce the lensing measurements in detail. Nevertheless, the best-fitting stellar masses are consistent within a few sigma between the C1 and C2 bins for all the runs and we therefore jointly fit them to derive a single stellar mass cut (for each feedback variant) for the kSZ effect CMASS analysis. Furthermore, we highlight that the kSZ effect measurements probe relatively small radii, in the 1-halo regime, where the simulation predictions match the CMASS lensing measurements well even for the C1 bin.
@@
[…]
We see visually from Fig. 2 and from the computed number of standard deviations that the fiducial calibrated FLAMINGO run is statistically ruled out by the kSZ effect measurements, at $\approx$7 sigma, for both the LOWZ and CMASS samples, which are independent. Only the strongest feedback model in this comparison (fgas-$8\sigma$) is formally consistent with the measurements, with a $\approx 1\sigma$ deviation from the measurements. Note that in computing the number of standard deviations we have not marginalised over uncertainties in cosmology (which we expect to be small, as discussed below) or uncertainties in the velocity reconstruction used in the observational measurements. Thus, the quoted level of tension may be slightly overestimated.

### v2

@@ Abstract
pubyear: 2024pagerange: FLAMINGO: combining kinetic SZ effect and galaxy-galaxy lensing measurements to gauge the impact of feedback on large-scale structureC
@@ 1 Introduction
Various possible solutions have been put forward to potentially reconcile the low-redshift LSS observations with the primary CMB and BAO data. These include unidentified or mischaracterised systematic uncertainties in the LSS observations, or possibly even in the primary CMB measurements. On the theoretical side, LSS tests of cosmology often probe deep into the non-linear regime, and it has been proposed that new physics on those scales could explain the tension (e.g., Amon & Efstathiou 2022; Preston et al. 2023), such as new dark sector models (e.g. Rogers et al., 2023; Elbers et al., 2025), or a mis-understanding of galaxy formation effects (e.g. McCarthy et al., 2018). Modelling non-linear scales necessitates cosmological simulations (or models calibrated on such simulations) to predict the clustering of matter on small scales and at late times. Additionally, as matter collapses to form self-gravitating halos, densities increase to the point where radiative cooling of the gas becomes efficient, leading to further collapse and galaxy formation. This process is accompanied by various energetic feedback mechanisms related to star formation and the accretion of matter onto supermassive black holes.
@@ 2.1 FLAMINGO simulations
The FLAMINGO suite consists of 16 hydrodynamical simulations presented in Schaye et al. (2023), two decaying dark matter variants in Elbers et al. (2025), plus two new hydrodynamical simulations introduced here (see below), and 12 gravity-only simulations. The suite has variations in resolution, box size, subgrid modelling, and cosmology. We mostly use the intermediate resolution ($m_{\mathrm{gas}}=1.09\times 10^{9}$ M) in box sizes of (1 Gpc)3. These simulations use $2\times 1800^{3}$ gas and dark matter particles and $1000^{3}$ neutrino particles, and most adopt cosmological parameters corresponding to the maximum likelihood DES Y3 ‘3×2pt + All Ext.’ $\Lambda$CDM cosmology (Abbott et al., 2022), which we refer to as ‘D3A’. These values assume a spatially flat universe and are based on the combination of constraints from DES Y3 ‘$3\times 2$-point’ correlation functions: cosmic shear, galaxy clustering, and galaxy-galaxy lensing, with constraints from external data from BAO, redshift-space distortions, SN Type Ia, and Planck observations of the CMB (including CMB lensing), Big Bang nucleosynthesis, and local measurements of the Hubble constant (see Abbott et al. 2022 for details). We also consider two alternative cosmologies: a run with the Planck 2018 maximum likelihood cosmology (‘Planck’; Planck Collaboration et al. 2020) and the ‘lensing cosmology’ from Amon et al. (2023) (‘LS8’). The latter model has a lower amplitude of the power spectrum, $S_{8}=0.766$, compared with 0.815 and 0.833 for D3A and Planck, respectively. We assess the resolution dependence of our results by comparing with a higher resolution run, labelled L1_m8. This run adopts the fiducial D3A cosmology in a (1 Gpc)3 volume but uses $2\times 3600^{3}$ gas and dark matter particles and $2000^{3}$ neutrino particles. The mass resolution is therefore a factor of 8 higher for this run (i.e., $m_{\mathrm{gas}}=1.34\times 10^{8}$ M) compared to the fiducial resolution runs.
The simulations were run with the cosmological smoothed particle hydrodynamics and gravity code SWIFT (Schaller et al., 2024b) using the SPHENIX SPH scheme (Borrow et al., 2022). The initial conditions are obtained from a modified version of monofonIC (Hahn et al., 2021; Elbers et al., 2022), and neutrinos are implemented with the $\delta f$ method (Elbers et al., 2021). The modelling of subgrid physics (star formation, stellar evolution, radiative cooling, and sources of feedback) is described in Schaye et al. (2023) and references therein.
@@ 2.1.1 Lightcones: HEALPix maps and galaxy/halo catalogs
The HEALPix map resolution is set to $N_{\text{side}}=16384$, which gives a maximum pixel radius of 13.46 arcseconds and $12*16384^{2}$ pixels in each full sky map. We note that the number of pixels exceeds the size of a signed 32-bit integer ($2^{31}$), which would prevent us from smoothing the kSZ effect maps with the ACT beam (necessary for a like-with-like comparison) using the Healpy smoothing function (sphtfunc.smoothing), as the function can currently only handle a maximum $N_{\text{side}}$ of 8192. We therefore downsample the kSZ effect maps to this resolution using the Healpy function pixelfunc.ud_grade, preserving the mean of the map in the downsampling operation. We describe the production of galaxy-galaxy lensing and kSZ effect maps in Sections 3.2 and 3.3, respectively.
Computational limitations prevent us from running a structure finder on-the-fly during the simulation. Instead, to produce galaxy/halo catalogs on the lightcone, structure finding [^1] We have checked that our results and conclusions are not strongly dependent on the choice of substructure finder, by comparing our fiducial results with those derived from employing an independent halo catalog derived using the VELOCIraptor package (Elahi et al., 2019). is done in post-processing on the snapshot particle data using a modified version of the HBT-HERONS algorithm (Han et al., 2012, 2018; Forouhar Moreno et al., 2025). Note that at $z\leq 3$ snapshots are also written out with a frequency of $\Delta z=0.05$, which was chosen to minimise any issues arising from the evolution of galaxy/halo properties between the snapshots and the HEALPix maps. We read in the snapshot subhalo catalog corresponding to a given snapshot and then read in a spherical shell from the black hole (BH) particle lightcone which spans the redshift range that is half way to the previous snapshot to halfway to the next snapshot. Every time a BH particle in the lightcone shell appears as the most bound BH in a subhalo in the snapshot (identified by matching their unique particle IDs), we place the subhalo at the BH particle’s position [^2] Note that subhaloes can potentially contain multiple black holes. To address this issue, we identify which of the BHs in a subhalo exist at both the next and previous snapshots and we pick the most bound BH particle from that subset to use as a tracer. In cases where the most bound BH survives, we then use the same BH as before. If the most bound BH disappears because it merges with a more massive BH before the next snapshot, we pick the next most bound BH in the same subhalo (which is not about to get swallowed) and use that as the tracer. in the halo lightcone. We repeat this process for every snapshot to make the full lightcone halo catalog. We use the halo lightcone to provide the locations of the selected galaxies for stacking the kSZ effect and galaxy-galaxy signals in the maps. We have verified that the profiles extracted from the lightcone-based HEALPix maps using positions from the constructed halo lightcone precisely match profiles constructed directly from the snapshots using the snapshot halo catalogs.
@@ 3.1 Galaxy selection
When selecting galaxies we use the true stellar mass predicted by the simulations. Observationally measured stellar masses have uncertainties, however, and an interesting question is whether our results would be impacted by folding in such uncertainties in the selection of our simulated galaxies. For the same reasons argued above, we do not expect such uncertainties to impact our results since the selection of galaxies (with or without observational uncertainties factored in) is forced to match the galaxy-galaxy lensing measurements. Nevertheless, we explicitly test this hypothesis below.
Note that an alternative possibility for selection would be to select systems based on their halo masses rather than their stellar masses, and to constrain the selected halo mass range based on fits to the galaxy-galaxy lensing measurements as described above. But such a selection would essentially correspond to a central-only population, since virtually all satellites would have halo masses well below the scale of interest here ($\sim 10^{13}\textrm{M}_{\odot}$) due to tidal stripping by their more massive host and would therefore either not be included in the selection or make no meaningful contribution to the stack. Selection by stellar mass, on the other hand, is closer to the real observational selection function in BOSS and allows us to naturally include both centrals and satellites in the selected simulation population.
@@ 3.2 Galaxy-galaxy lensing profiles
As noted previously, we downsample our full resolution $N_{\text{side}}=16384$ kSZ effect maps to $N_{\text{side}}=8192$ so that we can then smooth the maps with ACT beam. Schaan et al. (2021) used coadded ACT DR5 (day+night) maps in their study. We therefore retrieve the corresponding measured ACT beam for this dataset from the NASA Lambda website [^4] https://lambda.gsfc.nasa.gov/product/act/actpol_dr5_coadd_maps_info.html and convolve it with the simulated maps in multipole space using the Healpy function sphtfunc.almxfl.
Apart from convolution with a realistic beam, our maps are idealised, in that they do not contain realistic noise. We further assume that the kSZ effect signal can be perfectly recovered in the observational measurements, which would appear to be a strong assumption in the face of significant foreground and background contaminant signals, including the primary CMB, the tSZ effect, radio sources, the CIB, and so forth. However, in general these sources of contamination are not expected to correlate with the velocity field of the selected galaxies and are therefore suppressed in the velocity stacking process (see discussion in Schaan et al. 2021). Nevertheless, some contamination may still arise from the sources that are truly correlated with the selected haloes (i.e., part of the 2-halo term) and it will be important to evaluate their potential effects by constructing realistic mocks from the hydro simulations that contain all relevant signals. We leave this for future work (T. Yang, in prep).
@@ 4.1 Dependence on feedback and cosmology
The solid curves show the best-fitting profiles for each of the FLAMINGO runs, which fall nearly on top of each other. The legend provides the best-fitting ($\log_{10}$) minimum stellar mass (in solar masses) of the simulated galaxy selection for each of the runs. For example, in the LOWZ-L1 bin, the FLAMINGO fiducial feedback run with a minimum stellar mass of $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.24\pm 0.03$ provides the best fit to the DES Y3 + KiDS 1000 $\Delta\Sigma$ measurements. The error bars reflect the $2\sigma$ uncertainties (95% confidence) on the best-fitting minimum mass given the uncertainties on the lensing measurements. The best-fitting minimum stellar mass and its uncertainties are estimated by first calculating the $\Delta\Sigma$ profiles at different values of the minimum stellar mass and then interpolating to obtain the best fit result. The bottom right panel of Fig. 1 shows the process, where the dotted curves correspond to the predicted $\Delta\Sigma$ profiles for the fiducial FLAMINGO run with $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]$ values ranging between 10.9 and 11.6 in steps of 0.1 dex. We then interpolate the $\Delta\Sigma$ values to a much finer grid of $\log_{10}{M_{\text{star}}}$ Specifically, we linearly interpolate $\log_{10}\Delta\Sigma$ in stellar mass bins of width 0.01 dex at projected radius, $R$. The best-fitting minimum stellar mass is determined through $\chi^{2}$ minimisation with respect to the DES Y3 + KiDS1000 measurements and their uncertainties from Amon et al. (2023). Note that Amon et al. (2023) quote diagonal uncertainties for their combined DES Y3 + KiDS 1000 measurements only; i.e., the radial bins are assumed to be uncorrelated. Calculation of the off-diagonal elements of covariance matrix for the combined dataset is non-trivial and beyond the scope of this work.
@@
Figure 2: Comparison of the 150 GHz Planck+ACT kSZ effect temperature profiles of BOSS LOWZ and CMASS galaxies from Schaan et al. (2021) (data points with $1\sigma$ error bars) with predictions from the fiducial FLAMINGO run (L1_m9) and the fgas feedback variants. The bottom x-axis shows the angular scale (in arcmins) while the top x-axis shows the comoving projected radius at $z=0.31$ and $z=0.54$ for the LOWZ and CMASS panels, respectively. The vertical dashed lines indicate the location of $R_{\textrm{500c}}$ derived from the mean halo mass of each redshift bin (see Section 4.3). The solid curves represent the predictions for the best-fitting minimum stellar mass (fitted to the lensing, see Fig. 1). The legend provides the number of standard deviations that the model deviates from the observational measurements with the error bars corresponding to the propagated uncertainties on the stellar mass cut given the uncertainties on the galaxy-galaxy lensing measurements. Note that the number of standard deviations is computed using the full covariance matrices of the data, which is important since the outermost bins are highly correlated. The error bars on the data points reflect only the diagonal elements of the covariance matrices. The bottom sub-panels in each of the main panels show the ratio of the fgas feedback variants and the observational data with respect to the fiducial model. For reference, the dotted curves in the right panel show the predicted $\Delta T_{\rm kSZ}$ profiles for the fiducial FLAMINGO run with minimum stellar mass $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]$ values ranging between 10.9 and 11.6 in steps of 0.1. The fiducial calibrated FLAMINGO run is statistically ruled out by the kSZ effect measurements at about the $5\sigma$ level, for both the LOWZ and CMASS samples which are independent. Only the two strongest feedback models in this comparison (fgas-$4\sigma$, fgas-$8\sigma$) are formally consistent with the measurements.
@@
Comparing the fits across the two LOWZ bins (top two panels), for the majority of the cases the preferred minimum stellar masses are consistent within a few sigma for a given run. For example, the fiducial FLAMINGO model prefers a minimum stellar mass of $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.24\pm 0.03$ for the LOWZ-L1 bin and $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.29\pm 0.02$ for the LOWZ-L2 bin. As the two LOWZ bins are consistent with a single LOWZ selection for a given model, we jointly fit the LOWZ L1 and L2 bins to determine the minimum stellar mass cut for the kSZ effect predictions. The same is true for the higher redshift CMASS bins (C1 and C2, with $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.14\pm 0.05$ and $\log_{10}[M_{\text{star}}/\textrm{M}_{\odot}]=11.23\pm 0.04$, respectively), although we note that for the CMASS-C1 bin the shape of the best-fitting simulated profiles do not match the measurements perfectly. In particular, we note that the 1-halo regime dominates the fit (in terms of signal to noise), and the simulations provide a good match to the data there, but they predict a signal that is too large compared to the measurements at larger radii, in the 2-halo regime. The CMASS-C1 bin was also identified as an outlier in Amon et al. (2023), who found that even their flexible HOD framework (with 5 free parameters) was unable to reproduce the lensing measurements in detail. Nevertheless, the best-fitting stellar masses are consistent within a few sigma between the C1 and C2 bins for all the runs and we therefore jointly fit them to derive a single stellar mass cut (for each feedback variant) for the kSZ effect CMASS analysis. Furthermore, we highlight that the kSZ effect measurements mostly probe relatively small radii, in the 1-halo regime, where the simulation predictions match the CMASS lensing measurements well even for the C1 bin.
An alternative to approach to handling the two redshift bins in the LOWZ and CMASS samples of Amon et al. (2023) would be to predict the kSZ effect profiles for each bin given their respective best-fit minimum stellar masses and then to average the profiles (e.g., by inverse-variance weighting). In practice, we find that this gives nearly identical results to our default method of combining the bins. Indeed, even if we uniformly applied the lower or higher of the two stellar mass estimates to the whole sample, our general conclusions with regards to the strength of feedback required to match the kSZ effect measurements would be unchanged.
@@
We see visually from Fig. 2 and from the computed number of standard deviations that the fiducial calibrated FLAMINGO run is statistically ruled out by the kSZ effect measurements, at $\approx$5 sigma, for both the LOWZ and CMASS samples, which are independent. Only the two strongest feedback models in this comparison, fgas-$4\sigma$ and fgas-$8\sigma$, are formally consistent with the measurements, with $\approx 1.5\sigma$ and $\approx 0.5\sigma$ deviations from the measurements, respectively. Note that in computing the number of standard deviations we have not marginalised over uncertainties in cosmology (which we expect to be small, as discussed below) or uncertainties in the velocity reconstruction used in the observational measurements. Thus, the quoted level of tension may be slightly overestimated.
@@
[…]
As discussed in Section 3.1, observational stellar masses have measurement uncertainties and it is interesting to ask what the effect of such uncertainties might be if we applied them to the simulated galaxies and repeated the above analysis. Behroozi et al. (2019) find that a log-normal scatter with a standard deviation $\sigma(\log_{10}M_{\text{star}})=$min$(0.070+0.071z,0.3)$ dex describes typical random measurement errors in the observed stellar masses. As a test, we have applied this scatter to the true simulated stellar masses to mimic an observed stellar mass (note that we also applied this scatter during the calibration of FLAMINGO; see Schaye et al. 2023). We then analysed the simulations as described above, by determining the best-fit minimum stellar mass cut required to match the galaxy-galaxy lensing measurements and then predicting the kSZ profiles for this selection.
We find that the best-fit minimum stellar mass required to match the galaxy-galaxy lensing measurements typically increases by 0.04-0.05 dex with respect to our fiducial analysis with no measurement scatter. This reflects the fact there are more lower mass objects than high mass objects, thus leading to a slight net up-scattering. To reproduce the galaxy-galaxy lensing signal in the presence of lower stellar mass galaxies entering the selection, a slightly higher minimum stellar mass cut is therefore required to match the stacked lensing profiles, compared to the case with no scatter. Nevertheless, the best-fit mean halo mass is virtually identical to the no-scatter case, as is the predicted stacked kSZ profile. We therefore conclude that measurement uncertainties in the observed stellar masses do not significantly impact our results or conclusions, by virtue of the fact that the selection is constrained to match the lensing signal.
We have also tested the sensitivity of our results and conclusions to numerical resolution, by analysing the high-resolution calibrated FLAMINGO model (L1_m8) and comparing the results with the fiducial resolution, calibrated model (L1_m9). This comparison is presented in Appendix B. In short we find that the high-resolution run displays a similar level of tension with the observed kSZ effect measurements of Schaan et al. (2021) compared to the fiducial resolution, suggesting that our conclusions are robust to changes in resolution.